# dsld: A Socially Relevant Tool for Teaching Statistics

### v1

@@ Abstract
The growing power of data science can play a crucial role in addressing social discrimination, necessitating nuanced understanding and effective mitigation strategies of potential biases. “Data Science Looks At Discrimination” (dsld) is an R and Python package designed to provide users with a comprehensive toolkit of statistical and graphical methods for assessing possible discrimination related to protected groups, such as race, gender, and age. Our software offers techniques for discrimination analysis by identifying and mitigating confounding variables, along with methods for reducing bias in predictive models.
In educational settings, dsld offers instructors powerful tools to teach important statistical principles through motivating real-world examples of discrimination analysis. The inclusion of an 80-page Quarto book further supports users—from statistics educators to legal professionals—in effectively applying these analytical tools to real-world scenarios.
Keywords: Data science education; Statistical learning tools; Algorithmic fairness; Discrimination analysis; Confounding effects; Quarto notebook
@@ 1 Introduction
  • Addressing bias in prediction: This section focuses on the reduction of bias in the context of predictive algorithms. For example, consider a tool to aid in granting loan applications. If an applicant’s race is included as a predictor – either explicitly or through proxy variables – how can one mitigate its effect?
The value of the package is greatly enhanced via the use of a tightly integrated open source textbook, written in Quarto (Allaire et al., 2024). The book is not a user manual for the package, but instead serves as a detailed treatment of the statistical concepts themselves, illustrated with dsld examples.
@@ 2 The dsld Package and Quarto Book
The dsld package was developed in 2023 by seven undergraduates at the University of California, Davis, working under the direction of Norman Matloff. It currently consists of 24 functions, plus an associated open source textbook. It is available on CRAN, and the latest version can be found at https://github.com/matloff/dsld.
@@ 2.1 Relation to qeML package
As noted, the package consists of both analytical and graphical/tabular routines. Many of the former functions are wrappers for routines in other packages, adding discrimination-specific functionality.
In the context of using dsld for educational purposes, simplicity of the interface is of high importance to ensure accessibility across a broad student audience. It attains this in many cases via its use of the qeML package. The latter is designed to provide an exceptionally simple, direct user interface. For instance, to fit a linear model to predict mpg in R’s built-in mtcars dataset, in qeML one makes the simple call:
@@
No preparatory calls, no setup of any kind, just an easy request for a model fit. To do the same, but using random forests, just call:
@@
Use qeKNN(data = mtcars, yName = ‘mpg’) for K-Nearest Neighbors, and similar commands for other methods. Flexibility is retained via the number of default arguments.
@@ 2.2 Introducing dsld
Just as all qeML argument lists begin with data and yName, the same is true with most dsld functions, with a standard third argument being sName.
Furthermore, we see that dsldLinear has some new arguments not present in qeML, which are central to the discrimination analysis:
@@
If interactions = TRUE is set, interactions between the sensitive variable and the predictors/features in our dataset will be modeled. Actually, this approach involves fitting separate linear models for each level of the sensitive variable, which is statistically equivalent to modeling these interactions directly. The presence or absence of interaction terms plays an important role in the package and in the Quarto book. Is this a substantial difference in mean Y, given the co-variates X, across the levels of the R factor S? If so, is this difference uniform with respect to X?
@@
The package includes numerous graphical/tabular functions, some of them wrappers to existing packages, such as freqparcoord, and others standalone. These are used both for preliminary exploration of one’s data, and also for visual illustrations of the results found analytically.
For example, one might use dsldLogit to estimate parameters in a logistic model predicting probability of passage of the bar examination (binary response), based on LSAT scores provided in on the Law School Admissions dataset. We can then supplement these results with graphics such as Figure 1. Suppose we restrict the analysis to students with low GPAs. At initial glance, there appears to be significant racial differences at lower LSAT values, but these differences begin to diminish at higher scores. Notably, the probability curve for Asians shows an unexpected decline at higher LSAT values, warranting additional investigation. It is important to note that these results are purely exploratory, and further analysis is required to draw more definitive conclusions.
Figure 1: Predicted probabilities of passing the bar exam based on LSAT scores, illustrating racial disparities among various groups. Results are limited to students with GPAs below 2.70.
@@ 2.3 The Role of Nonparametric Regression Models
One may address nonlinearity via low-degree polynomial models, but modern computing power is such that a laptop computer can easily handle full nonparametric regression methods, even in fairly large datasets. In dsld, we feature Random Forests (RFs) as our primary non-parametric approach (Hastie et al., 2009).
@@
We chose RFs for several reasons. First, it is one of the most widely recognized non-parametric regression methods among statisticians, developed primarily by statisticians in the mid-90s and 2000s (Breiman, 2001a). (Less well known, but also significant, is the work by computer scientists such as Tin Ho Kam and her coauthors (Ho, 1995).) Second, RFs can be easily explained to non-statisticians, due to their “if this, then that” flow chart nature.
@@ 3 Detecting Discrimination
Discrimination is a significant social issue in the United States and many other countries. There is lots of available data with which one might investigate possible discrimination. But how might such investigations be conducted? There is a wide range of classical parametric methods available for this purpose, and non-parametric regression methods have recently gained popularity, particularly in applications like HR management (Frissen et al., 2022). The dsld package offers both graphical and analytical tools to detect potential biases, which are particularly beneficial for students. These tools can enhance understanding and intuition during classroom discussions by connecting them to broader social contexts.
In this section, we will use the lsa dataset on law school admissions, focusing on race as our sensitive variable of interest. The racial categories included are Asian, Black, White, Hispanic, and Other. This dataset is available through the dsld package for further exploration and analysis.
@@ 3.1 Graphical/Tabular Methods
Effective graphs and visualizations can significantly aid in students’ understanding of data. The dsld package offers various graphical methods to investigate discrimination issues, including dsldDensityByS, dsldConditDisparity, dsldConfounders and dsldScatterplot3D.
Consider investigating potential discrimination in the context of college and graduate school admissions. There has been growing criticism of standardized testing, arguing it disproportionately favors students with higher family income and other resources (Foreiter, 2021). Indeed, many studies have suggested unfair discrepancies in test results between Black and White students (Dixon-Roman et al., 2013). As a result, many institutions have removed standardized testing requirements like the SAT and GRE from their application processes. This debate underscores the importance of examining biases and motivates our study of the lsa dataset.
@@
Suppose we want to analyze the effect of LSAT scores on passing the bar examination to investigate potential racial disparities. Similar to the example shown in Figure 1, our analysis now encompasses all students, including those with high GPAs. We can calculate the differences in the probability of passing the bar for a given LSAT score among various racial groups. The function dsldConditDisparity is particularly well-suited for this purpose, as it visualizes such conditional disparities and may provide insights into how LSAT scores and bar examination outcomes may differ among racial groups.
@@
Figure 2: Predicted probabilities of passing the bar exam based on LSAT scores, illustrating racial disparities among various groups. Results include all students in the lsa dataset.
@@
Note that the dsldConditDisparity includes an optional argument condits (“conditions”). One could restrict the domain of students to those with lower undergraduate GPAs ($\leq$ 2.70), as shown in Figure 1, to conduct similar analyses.
@@
To further investigate the relationship between race and LSAT scores, we can generate a density plot using the dsldDensityByS function. This method enables users to visualize densities of a response variable $Y$ segmented by a sensitive variable $S$, with bandwidth control. This visualization provides insight into potential differences of the distribution of LSAT scores across different racial groups.
This is complementary to our previous graph, Figure 2, which suggested that White students pass the bar exam at higher rates than other races, even at similar LSAT levels. This difference may be exacerbated if White candidates also tend to do better on the LSAT. Consider the following plot:
@@
Figure 3: Density plot of LSAT scores segmented by race. The bandwidth parameter functions similarly to the bin width in a histogram and allow adjustment of the plot’s granularity. Default bandwidth is set at 1.
The density plot in Figure 3 illustrates the distribution of LSAT scores across different racial groups, highlighting potential bias. However, that apparent effect may be influenced by other confounding variables. Confounders influence both dependent and independent variables, possibly leading to spurious associations (McNamee, 2003). The function dsldConfounders acts similarly, but it is Plotly-based, and thus features interactive user access.
@@
Analyzing relationships among LSAT scores, undergraduate GPA, and family income can reveal potential confounding interactions.We can visualize these relationships by race using dsldScatterPlot3D, which offers a 3-dimensional perspective on multivariable relationships and highlights potential disparities among racial groups in terms of family income and GPA.
@@
Users can also utilize methods offered by the dsld package to identify potential confounding variables in their own analyses. For instance, the dsldCHunting function employs random forests to evaluate the importance of predictors in predicting $Y$ (without $S$) and $S$ (without $Y$). These predictors are identified as “important predictors” of $Y$ and $S$ and aid in the detection of potential confounding factors. For each $i$ from 1 to intersectDepth, the function reports the intersection of the top $i$ important predictors of Y and $S$. This helps identify potential confounders, with larger values of $i$ indicating inclusion of progressively weaker predictors.
@@
In the context of the lsa dataset, importance measures provide insights into predictors influencing the probability of passing the bar, which may act as confounders concerning race. Table 1 above shows the importance measures: the first row indicates predictors for the probability of passing the bar from other parameters (excluding race), while the second row shows predictors for race from other variables (excluding bar passage). For instance, variables that predict $Y$ (probability of passing the bar) well include decile3, LSAT, and cluster. In predicting $S$ (race), the top predictors are again LSAT, decile3, and cluster. By focusing on variables correlated with both $Y$ and $S$, we can identify potential confounders. For example, the intersection of the top three predictors for $Y$ and $S$ includes decile3, LSAT, and cluster. The analyst has flexibility in choosing the number of confounding variables, ranging from just the top two to more, depending on the individual needs and objectives. Furthermore, results from dsldConfounders can further add to one’s analysis. In particular, this method provides a density plot similar to dsldDensityByS for continuous predictors. For categorical variables, it displays the frequency of the sensitive variable occurring at each level of that variable. The Quarto book offers additional insights of bias detection using various datasets and examples.
@@ 3.2 Analytical
As noted earlier, in a no-interactions model, we can speak of, for instance, the Black and White difference in mean LSAT scores, independent of the values of $X$. But if interactions of $X$ and $S$ are assumed, there is no notion of “the” Black vs. White difference. That difference will vary according to the values of $X$, so the function estimates the difference in mean $Y$ at values of $X$ that are of interest to the user, as specified in sComparisonPts.
In general, a linear model with interaction terms involving a categorical variable with $m$ levels amounts to fitting $m$ separate linear models.
@@
We see for example that the estimated difference in dummy variables for Black and White scores is 6.0000 points with a standard error of 0.1410. In view of the fact that LSAT scores in this dataset ranged from 11 to 48 points, a 6-point difference seems to be very concerning.
@@ 4 Addressing Bias in Predictive Algorithms
In predicting $Y$ from $X$, one may wish to avoid or limit the influence of $S$. For example, in evaluating an application for a mortgage, we hope that our assessment does not discriminate against women or minorities. Such issues have been the subject of intensive research in recent years. There is even a conference devoted to this effort, the ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT). Here, we present some of dsld’s capabilities in this regard.
The “Hello World” example for this area of research involves the COMPAS (Correctional Offender Management Profiling for Alternative Sanctions) algorithm developed by Northpointe, a commercial entity. The algorithm was designed to predict a defendant’s likelihood of recidivism, thus aiding judges in determining sentences. Note that, as a commercial product, the algorithm is a “black-box,” of completely unknown details. The algorithm has faced criticism following an investigation by the publication ProPublica, which suggested that the algorithm exhibited bias against Black defendants (Angwin et al., 2016). Northpointe has contested these findings, arguing that ProPublica’s analysis is flawed. ProPublica has issued a rejoinder, but in any case, the debate over COMPAS underscores the critical need to address fairness.
Beyond the legal system, predictive methods have rapidly expanded into various commercial applications in recent years, including cybersecurity, healthcare, and e-commerce (Sarker, 2021). These models play a pivotal role in decision-making processes with significant consumer impact. This rise has highlighted concerns about algorithmic fairness in predictive methods (Wehner and Köchling, 2020; Chen, 2023).
Broadly speaking, “fairness” in this context refers to minimizing the influence of sensitive attributes (e.g., race, gender, religion, etc.) on an algorithm’s predictions (Oneto and Chiappa, 2020).
There are several main issues:
  • Defining unfairness: Several measures have been proposed.
  • Reducing unfairness: For a given algorithm, how can we ameliorate its unfairness while maintaining an acceptable predictive accuracy level?
@@
The dsld package provides methods for fair predictions with an intuitive, user-friendly interface for various fairness-constrained modeling approaches. In this section, we will use the COMPAS and lsa datasets to the showcase tradeoff between fairness and accuracy on both regression and classification settings. For the COMPAS dataset, we will predict the probability of recidivism with race as the sensitive feature. For the lsa dataset, we will predict LSAT scores with race as the sensitive variable. Both datasets are accessible via the dsld package for further exploration and analysis.

### v2

@@ Abstract
The growing power of data science can play a crucial role in addressing social discrimination, necessitating nuanced understanding and effective mitigation strategies for biases. “Data Science Looks At Discrimination” (dsld) is an R and Python package designed to provide users with a comprehensive toolkit of statistical and graphical methods for assessing possible discrimination related to protected groups such as race, gender, and age. The package addresses critical issues by identifying and mitigating confounders and reducing bias against protected groups in prediction algorithms.
In educational settings, dsld offers instructors powerful tools to teach statistical principles through motivating real-world examples of discrimination analysis. The inclusion of an 80-page Quarto book further supports users—from statistics educators to legal professionals—in effectively applying these analytical tools to real-world scenarios.
Keywords: Data science; Fair machine learning; Statistics education; Discrimination analysis; Confounder analysis; Quarto notebook.
@@ 1 Introduction
  • Addressing bias in prediction: This section focuses on the reduction of bias in the application of predictive algorithms. For example, consider a tool to aid in granting loan applications. If an applicant’s race is included as a predictor – either explicitly or through proxy variables – how can one mitigate its effect?
The value of the package is greatly enhanced via the use of a tightly integrated open source textbook, written in Quarto (Allaire et al., 2024). The book is not a user manual for the package, but instead is a detailed treatment of the statistical concepts themselves, illustrated with dsld examples.
@@ 2 The dsld Package and Quarto Book
The dsld package was developed in 2023 by seven undergraduates at the University of California, Davis, working under the direction of Norman Matloff. It currently consists of 24 functions, plus an associated open source textbook. It is available on CRAN, and the latest version is always in https://github.com/matloff/dsld.
@@ 2.1 Relation to qeML package
As noted, the package consists of both analytical and graphical/tabular routines. Many of the former functions are wrappers for routines in other packages, adding discrimination-specifc functionality.
In the context of educational usage of dsld, simplicity of the interface is of high importance, in order to be accessible by a broad student audience. It attains this in many cases via its use of the qeML package. The latter is designed to provide an exceptionally simple, direct user interface. For instance, to fit a linear model to predict mpg in R’s built-in mtcars dataset, in qeML one makes the simple call:
@@
No preparatory calls, no setup of any kind, just an easy request for a fit. To do the same, but with random forests, just call:
@@
Use qeKNN(data = mtcars, yName = ‘mpg’) for K-Nearest Neighbors and so on. Flexibility is retained via the number of default arguments.
@@ 2.2 Introducing dsld
We see that dsldLinear has some new arguments not present in qeML, which are central to the discrimination analysis:
@@
If interactions is set, interactions between the sensitive variable and the predictors/features in our dataset will be modeled. Actually, separate linear models will be fit for each level of the sensitive variable, which is essentially statistically equivalent. The presence or absence of interaction terms plays an important role in the package and in the Quarto book. Is this a subtantial difference in mean Y, given the covariates X, across the levels of the R factor S? If so, is this difference uniform with respect to X?
@@
The package includes numerous graphical/tabular functions, some of them wrappers to existing packages such as freqparcoord and others standalone. These are used both for preliminary exploration of one’s data, and also for visual illustrations of the results found analytically.
For example, one might use dsldLogit to estimate parameters in a logistic model predicting passage of the bar examination, based on scores on the Law School Admission Test, and then supplement the results with graphics such as Figure 1. There seem to be substantial differences among the races for low LSAT values, but not as much at the higher end.
Figure 1: Predicted probability of passing the bar exam by LSAT scores highlighting racial disparities among different groups. Student with GPA below 2.70.
Just as all qeML argument lists begin with data and yName, the same is true with dsld, with a standard third argument being sName.
@@ 2.3 The Role of Nonparametric Regression Models
One may address nonlinearity via low-degree polynomial models, but modern computing power is such that a laptop computer can easily handle fully nonparametric regression methods, even in fairly large datasets. In dsld, we feature Random Forests (RFs) as our primary non-parametric approach (Hastie et al., 2009).
@@
We chose RFs for a couple of reasons. First, it is arguably the most familiar non-parametric regression method among statisticians, having been developed primarily in the mid-90s and 2000s by statisticians (Breiman, 2001a). (Less well known, but also significant, is the work by computer scientists such as Tin Ho Kam and her coauthors (Ho, 1995).) Second, RFs can be easily explained to nonstatisticians, due to their “if this, then that” flow chart nature.
@@ 3 Detecting Discrimination
Discrimination is a significant social issue in the United States and indeed in all countries. There is lots of available data with which one might investigate possible discrimination. But how might such investigations be conducted? There is a rich array of classical parametric methods for this, and recently nonparametric regression methods are in increasing use, such as in HR management (Frissen et al., 2022). The dsld package offers both graphical and analytical tools to detect potential biases, which are particularly beneficial for students. These tools can enhance understanding and intuition during classroom discussions by connecting them to broader social contexts.
In this section, we will use the lsa dataset on law school admissions, focusing on race as the sensitive variable. The racial categories included are Asian, Black, White, Hispanic, and Other. This dataset is available through the dsld package for further exploration and analysis.
@@ 3.1 Graphical/Tabular Methods
Effective graphs and visualizations significantly enhance students’ understanding of data. The dsld package offers various graphical methods to investigate discrimination issues, including dsldDensityByS, dsldConditDisparity, dsldConfounders and dsldScatterplot3D.
Consider investigating potential discrimination in the context of college and graduate school admissions. There has been growing criticism of standardized testing, arguing it disproportionately favors students with more family income and other resources (Foreiter, 2021). Indeed, many studies have suggested unfair discrepancies in test results between Black and White students (Dixon-Roman et al., 2013). As a result, many institutions have removed standardized testing requirements like the SAT and GRE from their application processes. This debate underscores the importance of examining biases and motivates our study of the lsa dataset.
@@
Suppose we wish to analyze the effect of LSAT scores on passing the bar examination to investigate potential racial disparities. We may calculate differences in the probability of passing the bar for a given LSAT score among different racial groups. The function dsldConditDisparity is particularly suited for this purpose, as it visualizes conditional disparities, providing insights into how LSAT scores and bar examination outcomes might vary between racial groups.
@@
Figure 2: Estimated probability of passing the bar exam by LSAT scores, highlighting racial disparities among different groups. All students.
@@
The function also includes optional argument condits (“conditions”). For instance, we could restrict the domain of students to those with lower undergraduate GPAs ($\leq$ 2.70) as shown in Figure 1 and conduct similar analyses.
@@
To further explore the relationship between race and LSAT scores, we can generate a density plot using the dsldDensityByS function. In general, this method enables users to visualize densities of a response variable $Y$ segmented by a sensitive variable $S$, with bandwidth control. This visualization provides insight into potential differences of the distribution of LSAT scores across different racial groups.
This is complementary to our previous graph, Figure 2, which indicated that Whites pass the bar exam at higher rates than do non-Whites, even at the same LSAT levels. That difference may be exacerbated if Whites also tend to do better on the LSAT. Let’s see:
@@
Figure 3: Density plot of LSAT scores segmented by race. The bandwidth parameter can be thought as analogous to controlling the bin width in a histogram to adjust the figure granularity. Default bandwidth is set at 1.
(The function dsldConfounders acts similarly, but it is Plotly-based, and thus features user interactive access.)
The density plot above illustrates the distribution of LSAT scores across different racial groups, highlighting a potential bias. However, that apparent effect may be influenced by confounding variables. Confounders influence both dependent and independent variables, possibly leading to spurious associations (McNamee, 2003).
@@
For example, analyzing relationships among LSAT scores, undergraduate GPA, and family income can reveal potential confounding interactions. We can visualize these relationships segmented by race using dsldScatterPlot3D using a 3D perspective, offering insights into correlations and disparities across racial groups with respect to family income and GPA.
@@
between race and LSAT scores that some had assumed. Non-White students appear to
consistently show lower scores across various income levels, prompting further
discussion on how race influences exam outcomes while in consideration of other
[…]
in R interactive mode for better visualization.
[…]
function to aid in visualizing data in more than two dimensions. The
[…]
columns, where each column corresponds to a vertical axis in the graph. For
every data row, a polygonal line is drawn horizontally, with its height at
vertical section \$i\$ representing the value of variable for that row. Typically,
each variable is centered and scaled. Each row of data generates a distinctive
[…]
\$m\$ most frequently occurring patterns from each level of the sensitive
variable \$S\$. Two patterns are considered equivalent if they are proximate in a
k-nearest neighbor sense, with \$k\$ being an argument passed by the user. The
following call form will create the plot displayed in Figure
@@
Users can utilize methods offered by the dsld package to identify potential confounding variables in their own analyses. For instance, the dsldCHunting function employs random forests to evaluate the importance of predictors in predicting $Y$ (without $S$) and $S$ (without $Y$). These predictors are identified as “important predictors” of $Y$ and $S$ and aid in the detection of potential confounding factors. For each $i$ from 1 to intersectDepth, the function reports the intersection of the top $i$ important predictors of Y and $S$. This helps identify potential confounders, with larger values of $i$ indicating inclusion of progressively weaker predictors.
@@
In the context of the lsa dataset, importance measures provide insights into predictors influencing the probability of passing the bar, which may act as confounders concerning race. Table 1 above shows the importance measures: the first row indicates predictors for the probability of passing the bar from other variables (excluding race), while the second row shows predictors for race from other variables (excluding bar passage). For instance, variables that predict $Y$ (probability of passing the bar) well include decile3, LSAT, and cluster. In predicting $S$ (race), the top predictors are again LSAT, decile3, and cluster. By focusing on variables correlated with both $Y$ and $S$, we can identify potential confounders. For example, the intersection of the top three predictors for $Y$ and $S$ includes decile3, LSAT, and cluster. The analyst has flexibility in choosing the number of confounding variables, ranging from just the top two to more, depending on the individual needs and objectives.
Results from dsldConfounders can further add to one’s analysis. In particular, this method provides a density plot similar to dsldDensityByS for continuous predictors. For categorical variables, it displays the frequency of the sensitive variable occurring at each level of that variable. The Quarto book offers additional insights of bias detection using various datasets and examples.
@@ 3.2 Analytical
As noted earlier, in a no-interactions model, we can speak of, for instance, “the” Black and White difference in mean LSAT scores, independent of the values of $X$. But if interactions of $X$ and $S$ are assumed, there is no notion of “the” Black vs. White difference. That difference will vary according to the values of $X$, so the function estimates the difference in mean $Y$ at values of $X$ that are of interest to the user, as specified in sComparisonPts.
In general, a linear model with interaction terms involving a categorical variable with $m$ levels essentially amounts to fitting $m$ separate linear models.
@@
We see for example that the estimated difference in dummy variables for Black and White scores is 6.0000 points with a standard error of 0.1410. In view of difference seems to be rather concerning.
@@ 4 Addressing Bias in Predictive Algorithms
Machine learning (ML) is becoming increasingly popular in statistics courses to help students analyze complex data patterns and create predictive models. Learning ML techniques, especially with considerations over social fairness, can help students understand how to build models that not only perform well but also address issues like bias and inequality in real-world applications.
The challenge of fair machine learning is as follows: in predicting $Y$ from $X$, we want to avoid the influence of $S$. In evaluating an application for a mortgage, we hope that our assessment will not discriminate against women or minorities. Such issues have been the subject of intensive research in recent years. There is even a conference dedicated to this effort, the ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT). In this section, we present a brief overview of fairness in ML and some of dsld’s capabilities in this regard.
The “Hello World” example for this realm of research involves the COMPAS (Correctional Offender Management Profiling for Alternative Sanctions) algorithm developed by Northpointe, a commercial entity. The algorithm was designed to predict a defendant’s likelihood of recidivism and aid judges in sentencing decisions. Note that, as a commercial product, the algorithm is a “black-box,” of completely unknown details.
The algorithm faced criticism following an investigation by the publication ProPublica, which claimed that the algorithm exhibited bias against Black defendants (Angwin et al., 2016). Northpointe contested these findings, arguing that ProPublica’s analysis was flawed. ProPublica issued a rejoinder, but in any case, the debate over COMPAS underscores the critical need to address algorithmic fairness.
In this context, “fairness” generally refers to minimizing the impact of sensitive attributes such as race, gender, or religion on an algorithm’s predictions (Oneto and Chiappa, 2020). There are several main issues:
  • Defining unfairness: How is unfairness in predictions defined? Several measures have been proposed.
  • Reducing unfairness: How can we ameliorate an algorithm’s unfairness while maintaining an acceptable level of accuracy? This is known as the Fairness-Utility tradeoff.
@@
The dsld package provides a convenient and intuitive interface to a range of different fairness-constrained modeling approaches.
We use the COMPAS and Law School Admissions datasets to demonstrate the tradeoff between fairness and accuracy in both regression and classification settings. For the COMPAS dataset, we predict the probability of recidivism using race as the sensitive feature, while for the Law School Admissions dataset, we predict LSAT scores with race as the sensitive variable. Both datasets are accessible via the dsld package for further exploration and analysis.
4.1 Measuring Fairness and Accuracy
One possible measure of fairness is to require that $\widehat{Y}$ and $S$ be statistically independent (Johndrow and Lum, 2019). However, this approach eliminates the use of proxies, which makes it difficult to achieve a Fairness-Utility Tradeoff. While many other fairness measures have been proposed, they can be restrictive and may not apply across different applications.
A broader view would be to minimize the correlation between $\widehat{Y}$ and $S$. See for instance: Li et al. (2023), Kozodoi et al. (2022), Deho et al. (2022), Mary et al. (2019), and Baharlouei et al. (2019) applies work from Lee et al. (2022) and Roh et al. (2023).
The dsld package continues this theme:
  • Measuring Fairness: Fairness can be measured by the correlation between $\widehat{Y}$ and $S$ using the Kendall Tau statistic. The aim is to minimize the absolute correlation between $S$ and $\widehat{Y}$.
  • Measuring Accuracy: In regression settings, we define accuracy as the mean absolute prediction error (MAPE). In classification, we use the overall misclassification rate. Students may also use additional measures such as F1 score, sensitivity, and specificity for further analysis.